OpenStudy (anonymous):

Solve r+-9=4 for r. (r plus negative nine equals four) A. -13 B. -5 C. 5 D. 13 Need help!! :/

OpenStudy (texaschic101):

r + (-9) = 4 --- distriubte the + through the parenthesis r - 9 = 4 --- add 9 to both sides r = 4 + 9 r = 13

OpenStudy (anonymous):

Thanks!!

OpenStudy (texaschic101):

anytime :)